OpenStudy (anonymous):

How do you solve? 3 1/3 - 2 2/5=

OpenStudy (sandra):

hmmm, well, first you have to convert into some common denominator

OpenStudy (sandra):

so first, let's convert the mixed numbers into straight fractions

OpenStudy (sandra):

3 1/3 = 9/3 + 1/3 = 10/3

OpenStudy (sandra):

2 2/5 = 10/5 + 2/5 = 12/5

OpenStudy (sandra):

so now we need a common denominator for 3 and 5

OpenStudy (sandra):

15 is one

OpenStudy (sandra):

so 10/3 - 12/5 = 10*5/3*5 - 12*3/5*3 = 50/15 - 36/15 = 14/15

OpenStudy (anonymous):

wait im still at 12/5 that still needs to be simplefied correct?

OpenStudy (sandra):

well, 12/5 can't be simplified, since they have no common divisor (a number that divides evenly into them besides one)

OpenStudy (sandra):

the point is, you can multiply anything by 1 and get an equivalent number

OpenStudy (sandra):

so if you have 10/3 and 12/5 , then to easily subtract one from the other you need to same number on the bottom for both fractions

OpenStudy (sandra):

the easiest way to do that is to find a common denominator - but that only works because you can multiply any number by 1 and you know you have the same number - but maybe in a different form (fraction)

OpenStudy (anonymous):

sorry im slow.. so im going to back up a little.. so now we have 9/3 - 12/5 and we are finding common denomiator for 3 & 5 correct?

OpenStudy (sandra):

so 10/3 - I can multiply this by 1/1, 2/2, 3/3, 4/4, 5/5 etc... and it will be the same number

OpenStudy (anonymous):

wich you said was 15

OpenStudy (sandra):

yes you're right, but 10/3

OpenStudy (sandra):

so 10/3, and 12/5

OpenStudy (sandra):

your thought process at this point should be - "what do both 3 and 5 divide evenly into - since then i can just multiply to get that number at the bottom"

OpenStudy (sandra):

so really you're looking at 3 and 5

OpenStudy (sandra):

what number do they commonly divide into?

OpenStudy (sandra):

the answer is 15 - and that's always a sure method to get a common denominator - just multiply them by each other

OpenStudy (anonymous):

ok and the next part is what gets me so we find their common denomitator what do we do witht the top numbers?

OpenStudy (sandra):

so now your thought process should be - ok - so I they both divide evenly into 15 - what version of "1" can I multiply both fractions by to get 15? since again - I know if I multiple *any* number by "1", it's equivalent - so it's really just another way to write the same number

OpenStudy (sandra):

so for 3 - I know I multiply it by 5 to get 15

OpenStudy (sandra):

so I multiple 10/4 x 5/5 (again, multiplying by 1 means I get the same thing, different way of writing it)

OpenStudy (sandra):

sorry 10/3

OpenStudy (sandra):

so 10/3 x 5/5 = 10x5/3x5 = 50/15

OpenStudy (anonymous):

I get what your saying but im confused... wich doesnt take much :(

OpenStudy (sandra):

now the second fraction, how do I turn 12/5 into (some number)/15

OpenStudy (sandra):

well, I can multiply the bottom by 3 to get 15 (the common denominator) - but to keep the make the top number equal I need to multiply the top by the same number

OpenStudy (sandra):

so 12/5 = 12/5 * 3/3 (1) = 12*3/5*3 = 36/15

OpenStudy (sandra):

and once you have everything with the same common denominator, you can subtract across

OpenStudy (sandra):

so 50/15 - 36/15 = 50-36/15 = 14/15

OpenStudy (sandra):

but let's practice common denominators

OpenStudy (sandra):

so let's say 2/3 and 3/6 are your fractions

OpenStudy (sandra):

what number do both 3 and 6 divide evenly into?

OpenStudy (anonymous):

common denominator would be 6? sorry 3

OpenStudy (sandra):

nope you were right the first time

OpenStudy (sandra):

6

OpenStudy (sandra):

since 6/3 = 3, and 6/6 = 1

OpenStudy (sandra):

so that's good =)

OpenStudy (sandra):

ok so now you want to think - how can I make the bottom number a six for both of them?

OpenStudy (sandra):

so 2/3 = ?/6

OpenStudy (sandra):

and the main thing to learn, is that you can multipel *anything* by 1 and get the same number

OpenStudy (sandra):

i.e. 1x1 = 1, 2x1=2, 3x1 = 3

OpenStudy (sandra):

but that works with fractions too

OpenStudy (sandra):

so for 2/3, if I want to make the bottom number a six, I know I multiply it by 2 -

OpenStudy (sandra):

so to CONVERT 2/3, to ?/6, I just multiply by 2/2 - since it's another way of writing 1 (x/x always equals 1)

OpenStudy (sandra):

so 2/3 = 2/3 * 2/2 = 4/6

OpenStudy (anonymous):

ok..ok

OpenStudy (sandra):

great =) . so now convert 4/8 to ?/16

OpenStudy (sandra):

your first thought should be - what do I multiply 8 by to get 16

OpenStudy (anonymous):

2

OpenStudy (sandra):

exactly =). so now you can multiply 4/8 x 2/2 (since again, you're just multplying by "1" when you multiply by 2/20 - and get the equivalent fraction with 16 on the bottom

OpenStudy (sandra):

so 4/8 *2/2 = 4*2/8*2 = 8/16

OpenStudy (sandra):

and check this out - you could get an infinite number of equivalent fractions by just multiplying by different forms of 1

OpenStudy (sandra):

so 2/3 = 2/3 * 1/1 = 2/3 (duh) 2/3 = 2/3*2/2 = 4/6 2/3 = 2/3*3/3 = 6/9 2/3 = 2/3*4/4 = 8/12 etc...

OpenStudy (sandra):

those are all equivalent fractions - i.e. 2/3 = 4/6 = 6/9 = 8/12

OpenStudy (sandra):

and that's all a common denominator wants you to to

OpenStudy (anonymous):

so earlier we just choose 2/2 becuase that was the common denominator?

OpenStudy (sandra):

well, we chose 2/2 because we knew it would get us an equivalent fraction - the same exact value

OpenStudy (sandra):

when finding a common denominator, you choose the bottom number so that it makes the same thing the other fraction can make if it's multiplied by some number

OpenStudy (sandra):

so e.g. 3/5 and 2/3

OpenStudy (sandra):

well, I can multiply the bottom number of the first one by something to get 15, and I can multiply the bottom of the second number by something to also make it 15

OpenStudy (sandra):

15 is my common denominator

OpenStudy (anonymous):

your so patient with me...

OpenStudy (sandra):

and to convert 3/5 to ?/15, I know I multiply the bottom by 3, so I know I have to muitiply the top by 3 too - since that guarantees me I get the same number, but just "over" 15 instead of "over" 5

OpenStudy (sandra):

so 3/5 = 3/5 *3/3 (3/3 is just "1") = (3*3)/(5*3) = 9/15

OpenStudy (sandra):

and it's ok =) fractions are HARD until you understand them

OpenStudy (sandra):

the whole point again is - you can freely convert fractions to different forms as long as you multiply by 1

OpenStudy (anonymous):

thanks... A LOT!

OpenStudy (anonymous):

:)

OpenStudy (sandra):

no problem =)

OpenStudy (sandra):

but let's do a couple more

OpenStudy (anonymous):

absolutely!

OpenStudy (sandra):

so let's say I want you to convert 1/4 into ?/16

OpenStudy (sandra):

what is 1/4 if you make it ?/16

OpenStudy (sandra):

so your first thought should be, what do I need to to multiply 4 by to get 16 - since those are the two bottom numbers?

OpenStudy (anonymous):

1/4 X 4/4?

OpenStudy (sandra):

exactly!

OpenStudy (sandra):

which is?

OpenStudy (anonymous):

4/16

OpenStudy (anonymous):

or 1/4?

OpenStudy (sandra):

yep =) now put it into your calculator to check

OpenStudy (sandra):

nope, you were right, 4/16

OpenStudy (sandra):

well, they are both right lol

OpenStudy (sandra):

the thing with fractions, is you can also divide the top and the bottom by the same number and get the same thing

OpenStudy (sandra):

so 4/16 = (4/4)/ (16/4) = 1/4 !

OpenStudy (anonymous):

didnt see that at all...i just simplefied

OpenStudy (sandra):

and you can go the other way too =)

OpenStudy (sandra):

e.g. you can get all of the equivalent fractions starting at 1/4, by multiplying by different forms of 1

OpenStudy (sandra):

so 1/4 = 1/4 x 1/1 = 1/4 (duh) 1/4 = 1/4 x 2/2 = 2/8 1/4 = 1/4 x 3/3 = 3/13 1/4 = 1/4 x 4/4 = 4/16 1/4 = 1/4 x 5/5 = 5/20 etc...

OpenStudy (sandra):

and yep, 4/16 is on that list

OpenStudy (sandra):

ok so, let's do 5/6 = ?/18

OpenStudy (anonymous):

5/6 x 3/3 ?

OpenStudy (sandra):

exactly =)

OpenStudy (anonymous):

15/18

OpenStudy (anonymous):

horray!!!